# A backdoor attack against link prediction tasks with graph neural networks

### v1

@@ 4.1 Attack overview
where $E_{ij}$ is the target node pair in graph G that are unlinked and $\widetilde{E_{ij}}$ denotes the target node pair after embedding the trigger. From equation 5, the first formula represents the effectiveness of the backdoor model, which means that the backdoor model $f_{b}$ predicts the link state of the original-unlinked target node pair $\widetilde{E_{ij}}$ embedded in the trigger to be the linked state $T$. The second formula represents the evasiveness of the backdoor model, which means that the backdoor model predicts the link state of the target node pair $E_{ij}$ without the trigger embedded with normal behavior and similar prediction accuracy as the clean model.
Figure 2: The proposed GNN backdoor attack framework. our backdoor attack consists of four steps: 1. The attacker gets the trigger node by analyzing the features of the nodes in the graph dataset. 2. The attacker selects the poisoned node pairs by a score function. 3. The attacker generates the poisoning graph by injecting the trigger into the above selected node pairs and change the link state between the above node pairs to be linked. And this dataset is used to train the GNN models to embed the backdoor in the models. 4. After the backdoor is embedded into the GNN models, the unlinked node pairs with the trigger node will activate the backdoor in the models, and they will be predicted as linked, while those without the trigger node will be predicted correctly.
@@
    Trigger generation: the features of the trigger node are generated by counting the information of the features of all nodes in the training data and selecting the set of features indexes with the least frequency of occurrence, as detailed in Section 4.2 below;
@@
    Poisoned node pairs selection: a scoring function is proposed to select suitable node pairs by scoring each node pair in the training graph, which will be embedded in the trigger node for poisoning, as detailed in Section 4.3 below;
@@
    Backdoor injection: inject the trigger node into the poisoned node pairs selected in step 2 above (i.e., the trigger node is linked to two end nodes of the poisoned node pairs, respectively), and mark the linking relationship between the original-unlinked target node pairs as linked. The training graph dataset after injecting the trigger is called the poisoned training dataset. This poisoned training dataset is then used to train the GNN models, which learns those linked relationship between the trigger node and poisoned node pairs and injects it as the backdoor into the models, as detailed in Section 4.4 below;
@@
    Backdoor activation: in the inference phase, when the trigger node is connected to the end nodes of unlinked target node pair respectively, the backdoor models will incorrectly predict that there is a link between these two nodes. And when the trigger is not present, the backdoor models work normally.
The backdoor attack method proposed in this paper has the following assumptions:
@@
    This backdoor attack targets the link prediction under transductive learning. The goal of this backdoor attack is to incorrectly predict node pairs that are unlinked as linked.
@@
    The attacker can access to the training data, including the topology and features of the nodes, and the attacker can add additional edges and nodes to the graph to generate data for poisoning.
@@
    The attribute matrix of the graph is binary.
4.2 Trigger Generation
In this paper, we use a single node as the trigger and generate the features of the trigger node by the following two steps:
1.Determine the indexes of the position where ”1” occurs least frequently in the features of all nodes of the training dataset through statistics;
zheng et al. in motif-Backdoor [14] found that using non-existent or infrequent substructures as the triggers are more effective in implementing backdoor attacks. Inspired by this method, we use a single node as the trigger, select the position indexes in the dataset where the features of all nodes has the least frequency of ”1”, and then set the features value of the corresponding position indexes of the trigger’s features to ”1” according to the indexes of this position, which can explicitly separate the trigger node from the normal nodes in terms of the feature dimensions, so as to form a strong association between the trigger node and the target link state in the backdoor models obtained in the training process.
Specifically, in an undirected unweighted attribute graph $G=(V,E,X)$, where $V=\{v_{1},v_{2},\cdots,v_{N}\}$ is the set of nodes, $E$ is the set of edges, $X=R^{N\times d},j\in\{1,2,\cdots,d\}$ represents the nodes attribute matrix, and $X_{:,j}$ represents the individual column vectors in the node attribute matrix. In order to determine the position where ”1” occurs least frequently in the nodes attribute matrix, we count the number of 1st order paradigms of each column vector of the above matrix $X$ as follows:
@@ 4.2 Trigger Generation
Since we only consider the case where the attributes of the nodes in the graph are binary, the above formula calculates the frequency of occurrence of ”1” in each dimension of the features of each node, which is denoted as $\{\alpha_{1},\alpha_{2},\cdots,\alpha_{d}\}$. We select the top $k$ with the smallest values, denoted as $\{\alpha_{m_{1}},\alpha_{m_{2}},\cdots,\alpha_{m_{k}}\}$, where the index set $M_{k}=\{m_{1},m_{2},\cdots,m_{k}\}\subseteq\{1,2,c\dots,d\}$ stands for the k positions with the smallest frequency of ”1” in features of all nodes.
Considering that in general graph datasets such as Cora, CiteSeer, etc., the average number of features contained in nodes is generally about 1% of the total number of feature dimensions, and if the trigger node carries a large number of features can be easily detected, so in order to make our trigger node more unnoticable, we generally set $k$ to be 1% of the total number of feature dimensions.
2. Modify the features value of the corresponding positions of the trigger node to ”1” based on the position indexes above.
At first, all the elements of the initial features of the trigger node are set to ”0”, and then according to the above $k$ position indexes, the elements of the corresponding positions of the features of the trigger node are modified to ”1”. We denote the trigger node by $n_{t}$ and its i-th feature value by $x_{t}^{i}$, the above operation can be expressed as
@@
After the above two steps, we have generated the trigger node, which will be used to poison the training dataset in the following.
4.3 Poisoned node pairs selection
@@
After generating the trigger node features, the next step is to select the suitable node pairs to inject the trigger node to generate effective poisoned training data. Based on the message-passing mechanism of GNNs, the selection of poisoned node pairs is crucial for backdoor attacks.
In a graph, connected nodes may have the same features [38, 39]. A previous work [18] has shown that the embedding of the trigger destroys the node similarity in the original feature space, which will reduce the efectiveness of GNNs learning representations and decrease the efciency of downstream tasks. So the random injection of the trigger node destroys the node similarity of the feature space , and when the trigger node links different nodes, the GNNs aggregate confusing information that further misleads it to make wrong decisions[18]. To reduce this effect, we want all nodes embedded with the trigger to have links between them, i.e., the poisoned nodes form a complete graph between them, which can poison fewer nodes while the number of poisoned node pairs is fixed, and also ensures that all nodes containing trigger nodes will have links between them, which will enhance the effectiveness of the GNNs for backdoor feature learning. Besides, in order to be more conducive to the implementation of backdoor attacks, in the process of injecting the trigger node, we tend to select node pairs with sparse node features (meaning that the number of non-zero elements of the node’s features is as small as possible) to inject the trigger node. Because if our trigger node is connected to a node pair with dense features, then the backdoor model will have difficulty in distinguishing the features of the trigger node from the features of the normal nodes during the training process, which may lead to a decrease in the effect of the backdoor attack. In order to select node pairs with sparse features for posioning, we define the following node pairs scoring function (NPS) to measure the suitability of the node pairs in the training data to be used for injecting the trigger. The NPS function is defined as follows:
$NPS(v_{i},v_{j})=\|X_{v_{i}}\|_{1}+\|X_{v_{j}}\|_{1}$ (8)
@@
where, $(v_{i},v_{j})$ is any unlinked node pair in the training data, $\|X_{v_{i}}\|_{1}$ and $\|X_{v_{j}}\|_{1}$ represent the number of non-zero elements of the feature vectors of nodes $v_{i}$ and $v_{j}$, respectively. As can be seen, the sparser the features of the node pairs, the smaller the NPS values.
@@
We compute and sort all the unlinked node pairs in the training data according to the above NPS function, and select $q$ unlinked node pairs with the smallest NPS values as the target node pairs to be injected with the trigger, where, $q=p\cdot|V|$, $p$ is the poisoning rate, which represents the ratio of the number of poisoned node pairs to the total number of nodes in the graph.
4.4 Backdoor injection
@@
After generating the trigger node and selecting the target node pairs for injecting the trigger, the next step is to inject the trigger node into the above selected poisoned node pairs and change the link state between them from unlinked to linked, thus obtaining the poisoned graph $\widetilde{G}$. Finally, by training the GNN models with the above poisoned training data, the backdoor is embedded in the target model. The overall algorithm flow is shown in algorithm 1.
4.5 Backdoor Activation
@@
In the inference phase, when the trigger node is connected to two end nodes of unlinked target node pair respectively, the backdoor models incorrectly predict that there is a link between these two nodes. And when the trigger is not present, the backdoor models work normally.
@@
0:  $G=(V,E,X)$,poisoning rate $p$ Calculate the occurrence frequency $\alpha_{j}$ of each feature by equation 6 Select the first k features with the lowest occurrence frequency to add to the index set $M_{k}$ Generate the trigger node $n_{t}$ by equation 7 Calculate the score of node pairs by equation 8 Select poisoned node pairs for poisoning by their scores Inject the trigger node into the poisoned node pairs Changes the link status of the poisoned node pairs to linked to generate poisoning graph $\widetilde{G}$ Train the model on poisoning graph $\widetilde{G}$ to obtain $f_{b}$
0:  backdoor model $f_{b}$ and trigger $n_{t}$
5 Experiment
To verify the effectiveness of the proposed backdoor attack method, we conduct the following experiments: initially, we test the attack success rates of the backdoor models generated under different poisoning rates on samples with the trigger and the prediction accuracy on clean samples to evaluate the attack effectiveness and stealthiness of backdoor models, as described in Section 5.2; secondly, we compare these results with the baseline to assess the effectiveness of the backdoor attack method proposed in this paper, as described in Section 5.3. Then we test the effects of different budgets $k$ on the attack success rates of the backdoor models and the prediction accuracy of the backdoor model on clean samples respectively, as described in Section 5.4. Finally we conducted the ablation studies to confirm the effectiveness of the scoring-based poisoned node pairs selection method used in the paper, as described in Section 5.5.
5.1 Experiment setting
5.1.1 Datasets
We evaluate our method in four graph benchmark datasets: Cora[40], CiteSeer[41], CS[42], and Physcics[42]. They are all undirected graphs composed of literature citation networks, where nodes represent papers, edges represent citation relations, node features are word vectors, and node labels are paper topics. Table 2 summarizes the basic statistics of these four graph databases.
Table 2: The dataset statistics
@@ 5.1.1 Datasets
5.1.2 Link prediction models
In order to evaluate the effectiveness of our proposed attack method, we use four popular GNN models on link prediction, that is, Graph Auto-Encoder (GAE)[7], Variational Graph Auto-Encoder (VGAE) [7], Adversarial Regularized Graph Auto-Encoder (ARGA) [8], and Adversarial Regularized Variational Graph Auto-Encoder (ARVGA)[8]. We briefly describe these models as follows:
@@ 5.1.2 Link prediction models
5.1.3 Parameter setting
We construct GAE, VGAE, ARGA and ARVGA models for link prediction. For the parameters of the link prediction models, we retained the settings described in the corresponding papers. That is, For VGAE and GAE, we train for 200 iterations using Adam with a learning rate of 0.01. We use a 32-dim hidden layer and 16-dim latent variables in all. For ARGA and ARVGA, we train all auto-encoder related models for 200 iterations and optimize them with the Adam algorithm. Both learning rate and discriminator learning rate are set for the ARGA and ARVGA. rate and discriminator learning rate are set as 0.001. We construct encoders with a 32-neuron hidden layer and a 16-neuron embedding layer for all the experiments and all the discriminators are built with two hidden layers (16-neuron, 64-neuron respectively). In our experiments, we repeat the above steps five times and averaged the results of the experiments.
5.1.4 Metrics
To evaluate the effectiveness and evasiveness of the backdoor attack method proposed in this paper, we use the following evaluation metrics.
@@ 5.1.4 Metrics
    Attack success rate (ASR): Attack success rate (ASR) represents the percentage of node pairs that are unlinked embedded in the trigger that the backdoor model predicts to be in a linked state. It is as follows:
    $ASR=\frac{Number\ successful\ attack\ node\ pairs}{Total\ number\ of\ attack\ node\ pairs}$ (9)
@@
    Benign performance drop (BPD): To evaluate the evasiveness of our attack method,we utilize the area under the curve,AUC[43], to represent the accuracy of the link prediction model.If, in $n$ independent comparisons,the link that exists gets a higher score than the link that does not exist $n^{{}^{\prime}}$ times, and they get the same score $n^{"}$ times, then AUC is defined as
    $AUC=\frac{n^{{}^{\prime}}+n^{"}}{n}$ (10)
    We then use the Benign Performance Difference (BPD), which measures the difference in prediction accuracy between the clean GNN model and the backdoor GNN model on clean graph. The BPD is defined as
@@
    where $AUC_{c}$ and $AUC_{b}$ represent the accuracy of clean model and backdoor model respectively.Lower BPD represents better performance, which indicates that backdoor model and clean model are closer to each other in terms of accuracy on clean samples, and the backdoor attack have better stealth.
@@
    Poisoning rate (p):The poisoning rate $p$ represents the ratio of the number of target node pairs poisoned to the total number of nodes in the graph. The lower the poisoning rate, the easier and more stealthier the backdoor attack will be.
@@
    Budget(k): The budget $k$ respresents the number of modified trigger features. budget is defined as follows:
@@
    where $k$ is the budget, $d$ is the total number of node feature dimensions, and $\lambda$ represents the number of modified trigger node features as a percentage of the total number of node feature dimensions.
5.1.5 Baseline
In order to evaluate our attack method, we compare the only current paper on link prediction for static graph data, LB[22]as baseline, with our method.
5.2 The effects of different poisoning rates
We test the ASR and BPD of the backdoor models generated against the four target models GAE, VGAE, ARGA and ARVGA with different poisoning rates on four datasets to evaluate the impact of the poisoning rate on the effectiveness and evasiveness of the backdoor attack method proposed in this paper, where the poisoning rates in each dataset of the experiments are in order of 0.2%,0.5%,1%,2%, 5%. The results of the experiments are shown in the table 3.
Table 3: Impact of poisoning rate on ASR and BPD.
Datasets p ASR BPD

### v2

@@ IV-A Attack overview
where $E_{ij}$ is the the target unlinked node pair without the trigger in graph G and $\widetilde{E_{ij}}$ denotes the target unlinked node pair injected with the trigger. The first formula of equation 5 represents the effectiveness of the backdoor attack, which means that the backdoored GNNs model $f_{b}$ incorrectly predicts link state of the target unlink node pair as the linked state $T$ due to the appearance of the trigger. The second formula of equation 5 represents the evasiveness of the backdoor attack, which means that the backdoored GNN model works well for the target unlinked node pair $E_{ij}$ without the trigger and its predicting accuracy is as close as possible to that of the corresponding clean model.
Figure 2: The proposed backdoor attack framework. The backdoor attack consists of four steps:1. The attacker generates the trigger node, as shown in red dot in the figure, the features of which are got by selecting the top k features with the least occurrence frequency, based on statistical information of the features of all nodes in the original training dataset. 2. The attacker selects unlinked node pairs for poisoning from the training graph through a score function, as shown in the figure, node A and node B are one of such selected unlinked node pairs. 3. The attacker generates the poisoning training dataset by injecting the trigger into the above selected node pairs (i.e. connect the trigger node to two end nodes of the node pairs, for example, connect the red trigger node to node A and node B in the figure) and change the unlinked state of the selected node pairs to linked state. Then the poisoning training dataset is used to train the GNN models to embed the backdoor in the models, which is called backdoored GNNs models. 4. In the inference phase of the backdoored GNN models, if the unlinked node pairs are connected to the trigger node , such as the node E and the node F in the sample with the trigger in the figure, they will activate the backdoor in the models and will be predicted incorrectly to be linked, while for the unlinked node pairs without the trigger node, such as the node C and node D in the benign sample in the figure, they will be predicted correctly to be unlinked.
@@
    Trigger generation: the features of the trigger node are generated by selecting the top k features with the least occurrence frequency, based on statistical information of the features of all nodes in the training data, as detailed in Section IV-B below;
@@
    Selection of node pairs for poisoning: a scoring function is proposed to select suitable node pairs in the training graph, which will be injected with the trigger node in the next step to generated poisoning training dataset. as detailed in Section IV-C below;
@@
    Backdoor injection: The poisoned training dataset is generated by injecting the trigger node into the selected node pairs from step 2 above and changing the link state of the selected unlinked node pairs to be linked. The poisoned training dataset is then used to train the GNN models to inject the backdoor, as detailed in Section IV-D below;
@@
    Backdoor activation: in the inference phase, when the trigger node is connected to the end nodes of target unlinked node pair respectively, the backdoor in the backdoored model will be activated and it will incorrectly predict that there is a link between these two unlinked nodes. And when the trigger is not present, the backdoored model work well.
The backdoor attack proposed in this paper has the following assumptions:
@@
    We focus on link prediction under transduction learning. The goal of this backdoor attack is to incorrectly predict unlinked node pairs as linked.
@@
    The attacker can access to the training data, including the topology and features of the nodes, and the attacker can add additional edges and nodes to the graph to generate poisoning data.
@@
    The feature matrix of the graph is binary.
IV-B Trigger Generation
In this paper, we use a single node as the trigger and generate its feature vector by the following two steps:
1.Identify top k indexes of dimensions where element “1” occurs least frequently in the binary feature vectors of all nodes in the training dataset.
Zheng et al. found in their research on motif-Backdoor [14] that using non-existent or infrequent substructures as the triggers is more effective in implementing backdoor attacks for GNNs. Inspired by this method, we use a single node as the trigger and distinguish the trigger node from other normal nodes in the training dataset in terms of the feature vector, so as to form a strong association between the trigger node and the linked state in the training process.
Specifically, in an undirected unweighted feature graph $G=(V,E,X)$, where $V=\{v_{1},v_{2},\cdots,v_{N}\}$ is the set of nodes, $E$ is the set of edges, $X=R^{N\times d},j\in\{1,2,\cdots,d\}$ represents features matrix of the nodes, the rows of which are d-dimensional feature vectors of the corresponding nodes. in order to determine the indexes of dimension where element “1” occurs least frequently in the binary feature vectors of all nodes in the training dataset. we compute the $L_{1}$ norm of each column vector of the above matrix $X$ as follows:
@@ IV-B Trigger Generation
where $X_{:,j}$ represents the d-dimensional column vectors in the node feature matrix. Since we only consider the case where the features of the nodes in the graph are binary, the above formula calculates the frequency of occurrence of ”1” in each dimension of the feature vectors of all nodes, which is denoted as $\{\alpha_{1},\alpha_{2},\cdots,\alpha_{d}\}$. We fututer select the smallest $k$ values among them, denoted as $\{\alpha_{m_{1}},\alpha_{m_{2}},\cdots,\alpha_{m_{k}}\}$, where subscripts set $M_{k}=\{m_{1},m_{2},\cdots,m_{k}\}\subseteq\{1,2,3\dots,d\}$ stands for the top $k$ indexes of dimension where element “1” occurs least frequently in the binary feature vectors of all nodes in the training dataset. $k$ is called trigger size which is the number of modified elements in the feature vector of the trigger node ,as described in the following step.
2. Modify the value of the corresponding dimension in the feature vector of the trigger node to ”1” based on the above dimension indexes $M_{k}$.
At first, all the elements in the feature vector of the trigger node are initially set to ”0”, and then according to the above $k$ dimension indexes, the value of the corresponding dimension in the feature vector of the trigger node are modified to ”1” . We denote the trigger node by $n_{t}$ and its i-th feature value by $x_{t}^{i}$, the above operation can be expressed as:
@@
After the above two steps, we have generated the trigger node, which will be used to poison the training dataset in the following step.
IV-C Selection of node pairs for poisoning
@@
After generating the trigger node features, the next step is to select the suitable node pairs to inject the trigger node to generate effective poisoning training data.
Considering that the poisoned training dataset is generated by connecting the trigger node with the end nodes of the selected unlinked node pairs in the original training graph. The links between the trigger node and the end nodes of these unlinked node pairs may affect learning the features of the trigger node by the GNN model. Due to the difficulty of GNNs in extracting hidden features from sparse node features (sparse meaning that the number of elements with value 1 in the node’s feature vector is far less than that of elements with value 0)[38, 39], these node pairs with sparse node features will have less impact on learning the features of the trigger node by the GNN model . Therefore, we select node pairs with sparse features to inject the trigger node so as to enhance the influence of the trigger node in the backdoor attack and form a strong association between the trigger node and linked state in the training process. The scoring function to select unlinked node pairs with sparse features to inject the trigger node is defined as follows:
$Score(v_{i},v_{j})=\|X_{v_{i}}\|_{1}+\|X_{v_{j}}\|_{1}$ (8)
@@
where, $(v_{i},v_{j})$ is an unlinked node pair in the training data, $\|X_{v_{i}}\|_{1}$ and $\|X_{v_{j}}\|_{1}$ represent the number of non-zero elements of the feature vectors of nodes $v_{i}$ and $v_{j}$, respectively. As can be seen, the sparser the features of the node pairs, the smaller the scores.
@@
For all the unlinked node pairs in the training data, we compute their value with the above score function respectively, and select $q$ unlinked node pairs with the smallest score values as the target node pairs to inject the trigger, where, $q=p\cdot|V|$, $p$ is the poisoning rate, which represents the ratio of the number of poisoned node pairs to the total number of nodes in the training graph.
IV-D Backdoor injection
@@
After generating the trigger node and selecting the node pairs for poisoning, the next step is to generate poisoning training dataset $\widetilde{G}$ by injecting the trigger node into the selected node pairs(i.e. connect the trigger node with two end nodes of the node pairs) and changing the unlinked state of the node pairs to be linked. Then the poisoning training dataset is used to train the GNN models to embed the backdoor in the models. The overall algorithm is shown in algorithm 1.
IV-E Backdoor Activation
@@
In the inference phase, if the trigger node is connected to two end nodes of unlinked target node pair respectively, the backdoor in the backdoored models will be activated and they will incorrectly predict that there is a link in the unlinked node pair, while for unlinked node pairs without the trigger node, the backdoored models work normally.
@@
0:  $G=(V,E,X)$,poisoning rate $p$
1:  for $i=1$ to $N$ do
2:     Calculate $\alpha_{i}$ of each feature dimension by equation 6
3:  end for
4:  Get the dimension indexes $M_{k}$ by selecting the k dimensions with minimum $\alpha_{i}$
5:  Generate the features $x_{t}$ of the trigger node by equation 7
6:  Calculate the score of node pairs by equation 8
7:  select unlinked node pairs for poisoning according the above score
8:  Generate the poisoning graph by injecting the trigger node into the selected unlinked node pairs and changing the link state of the unlinked node pairs to be linked
9:  Train the model on poisoning graph $\widetilde{G}$ to obtain $f_{b}$
9:  backdoored model $f_{b}$ and trigger $n_{t}$
V Experiment
The following four experiments have been conducted to evaluate the effectiveness of the proposed backdoor attack. Firstly, we generated the backdoored GNN models with different poisoning rates and tested the attack success rate of these backdoored models on samples with trigger as well as their prediction accuracy on clean samples respectively to evaluate the effectiveness and stealthiness of the proposed backdoor attack. as described in Section V-B. Secondly, we compared these results with the baseline to assess the performance of the proposed backdoor attack, as described in Section V-C. Next we tested the impacts of different trigger sizes on the attack success rates of the backdoored models and their prediction accuracy on clean samples respectively, as described in Section V-D. Finally we conducted the ablation studies to confirm the effectiveness of the score function used to select node pairs for poisoning on the performance of the proposed backdoor attack, as described in Section V-E.
V-A Experiment setting
V-A1 Datasets
We evaluate the proposed backdoor attack in four graph benchmark datasets: Cora[40], CiteSeer[41], CS[42], and Physcics[42]. They are all undirected graphs composed of literature citation networks, where nodes represent papers, edges represent citation relations, node features are word vectors, and node labels are paper topics. Table II summarizes the basic statistics of these four graph databases.
Table II: The dataset statistics
@@ V-A1 Datasets
V-A2 Link prediction models
In order to evaluate the effectiveness of the proposed backdoor attack, we use four popular GNN models for link prediction, that is, Graph Auto-Encoder (GAE)[7], Variational Graph Auto-Encoder (VGAE) [7], Adversarial Regularized Graph Auto-Encoder (ARGA) [8], and Adversarial Regularized Variational Graph Auto-Encoder (ARVGA)[8]. We briefly describe these models as follows:
@@ V-A2 Link prediction models
V-A3 Parameter setting
We construct GAE, VGAE, ARGA and ARVGA models for link prediction and adopt the same settings as those in the corresponding original papers. We construct two-layer GCNs with a 32-dim hidden layers and 16-dim hidden layers in all. For ARGA and ARVGA, the discriminators are built with two hidden layers (16-neuron, 64-neuron respectively). In the experiment, we repeated the experiment five times and averaged the results.
V-A4 Metrics
To evaluate the effectiveness and evasiveness of the backdoor attack proposed in this paper, we use the following metrics.
@@ V-A4 Metrics
    Attack success rate (ASR) refers to the ratio of the number of the attack node pairs (i.e. target unlinked node pairs with the trigger) predicted to be linked to the total number of the attack node pairs. It is as follows
    $ASR=\frac{N_{suc}}{N_{att}}$ (9)
    where $N_{suc}$ is the number of the attack node pairs predicted to be linked, $N_{att}$ is the total number of attack node pairs.
@@
    Benign performance drop (BPD): To evaluate the evasiveness of our attack, we utilize the area under the curve, AUC[43], to represent the accuracy of the link prediction model. If, among $n$ independent comparisons, the number of times that the existing link gets a higher score than the nonexistent link is $n^{{}^{\prime}}$, and the number of times they get the same score is $n^{\prime\prime}$, then the AUC is defined as
    $AUC=\frac{n^{{}^{\prime}}+0.5n^{{}^{\prime\prime}}}{n}$ (10)
    We then use the Benign Performance Difference (BPD), which measures the difference in prediction accuracy between the clean GNN model and the backdoored GNN model on clean graph. The BPD is defined as
@@
    where $AUC_{c}$ and $AUC_{b}$ represent the accuracy of clean model and backdoored model respectively. Lower BPD represents better performance, which indicates that backdoored model and clean model are closer to each other in terms of accuracy on clean samples, and the backdoor attack achieves more stealthy.
@@
    Poisoning rate (p): The poisoning rate $p$ represents the ratio of the number of target unlinked node pairs embedded with the trigger node to the total number of nodes in the graph. The lower the poisoning rate, the easier and more stealthier the backdoor attack will be.
@@
    Trigger size(k): The $k$ represents the number of elements in the feature vector of the trigger node that have been changed from an initial value of 0 to a value of 1 in the process of generating the trigger node. The trigger size is defined as follows
@@
    where $k$ is the trigger size, and $d$ is the total number of the trigger node feature dimensions, and $\lambda$ is the budget which represents the percentage of the modified features among all features in the trigger node.
V-A5 Baseline
In order to evaluate the proposed backdoor attack, we compare it with the only one state-of-the-art baseline LB[20] which is a backdoor attack method for link prediction tasks on static graphs, using subgraphs as triggers.
V-B Testing the impact of different poisoning rates on ASR and BPD
We test the ASR and BPD of the backdoored models generated from the four GNN models GAE, VGAE, ARGA and ARVGA with different poisoning rates on four datasets to evaluate the impact of the poisoning rate on the effectiveness and evasiveness of the backdoor attack method proposed in this paper, where the poisoning rates in each dataset of the experiments are 0.2%, 0.5%, 1%, 2%, 5%, respectively. The results of the experiments are shown in the table III.
Table III: The impact of the poisoning rate on ASR and BPD.
Datasets p(%) ASR(%) BPD(%)